# “Super-resolution” holographic optical tweezers array

### v1

@@ Results
Spot uniformity and light utilisation efficiency. Considering the scope of this study, the evaluation of the reproduced images should be based on the characteristic properties of light-spot arrays. One might assume that the FOI itself serves as a measure of the quality of the reproduced images. However, FOI primarily reflects the total correspondence between the output and target patterns. Therefore, we investigated the properties of the reproduced images in terms of the uniformity of the light spot array and light utilisation efficiency.
The spot uniformity and utilisation efficiency of the images reproduced using FOI- and RSS-designed CGHs are presented in Table 1. The results reveal that the FOI method exhibits small variations in spot power but demonstrates low light utilisation efficiency, despite employing phase-only modulation. In contrast, the RSS results display the opposite behaviour. Uniformity was suppressed in both numerical and experimental results, whereas the numerical and experimental values of light utilisation efficiency were similar. Thus, the FOI method can generate high-quality spot arrays in terms of uniformity and stability, although at the expense of efficiency.
@@
Figure 3: Spacing and separation of adjacent light spots. a, Focal pattern (top) and profile (bottom) of Airy disk [NA = 0.75(air)]. Red line in the bottom plot shows cross-sectional profile corresponding to the focal pattern, where $r_{\text{A}}$ refers to the Airy-disk radius. b, Focal patterns and profiles of incoherent sums of two Airy disks aligned at distances of $d=r_{\text{A}}$ (left) and $2r_{\text{A}}$ (right). In the bottom plots, thick green lines display the profiles of the incoherent sums, whereas thin red lines indicate those of individual Airy disks. Definitions of peak (P) and valley (V) are illustrated as well. c, (Solid green line)  The relationship between the VP ratio and spot spacing for incoherent sums of adjacent Airy disks is depicted, where the traditional Rayleigh diffraction limit corresponds to $\text{VP}=0.66$ at $d/r_{\text{A}}=1$. Open circles represent VP ratios for numerically reproduced images from FOI- (red open circles) and RSS-designed (blue open circles) CGHs, while closed circles represent those for the corresponding observed patterns, i.e., Figs. 2c (red closed circle) and d (blue closed circle), respectively. Error bars (SEs) are determined similarly to those in Table 1 but are smaller than the size of the plot symbols, except for the experimental result of an RSS-designed CGH.
@@
Relation between spacing and separation of spot array. FOI-designed holograms can generate high-quality spot arrays with extremely narrow spacing. To evaluate the behaviour of spot arrays within the diffraction-limit regions, the focal patterns calculated under a high-NA [= 0.75 (air)] condition are displayed in Fig. 3. Note that all numerical results in Fig. 3 were calculated using vector Debye formulas [26, 27, 28, 29, 30], as the focal patterns of coherent light are affected by the vector nature of the light amplitude under NA conditions greater than 0.70. The Airy disk pattern calculated assuming circularly polarised light is illustrated in Figure 3a.
@@
The Rayleigh-limit conditions portrayed on the left in Fig. 3b present the adjacent light spots (Airy disks) as inadequately separated, while well-separated light spots, as displayed on the right of Fig. 3b, are desirable for maintaining a high aspect ratio in applications such as laser processing and suppression of background noise in scanning laser microscopy. Therefore, a factor other than spot spacing is required for a proper evaluation of spot-array patterns while considering their applications.
@@
The valley-to-peak (VP) ratio of a spot array along with the inverse of the peak-to-valley ratio was introduced as a metric for spot separation. A smaller VP ratio indicates a clear separation of adjacent light spots. The VP ratio is advantageous because of its availability and robustness in derivation under unfavourable conditions compared to other metrics, such as the size of each light spot. To derive the VP ratio of the reproduced spot-array image, the peak positions were determined by fitting the reproduced images (Methods). The peak value refers to the mean of $I_{i}$ at the peak positions of all the bright spots, whereas the valley value refers to the mean of $I_{i}$ at all the midpoints between the adjacent peaks. Thus, the VP ratio was evaluated as the ratio of the valley value to the peak value.
@@
The blue (red) open circles displayed in Fig. 3c represent the VP ratios of numerically reproduced images from hologram patterns designed by the FOI (RSS) cost method, whereas the blue (red) closed circles indicate those derived from the experimental images. In terms of the VP ratio, the proposed FOI method outperforms the conventional RSS method in the range of $(1<d/r_{\text{A}}<2)$.
Discussions
@@
The super-resolution criterion can be defined as both definite and quantitative. In considering the relationship between physical quantities involving spot separation, the traditional Rayleigh diffraction limit is established for an incoherent superposition of two Airy disks. Thus, this approach should be logically followed. We propose to determine whether a spot array exhibits “super-resolution” by examining the relationship between the spacing $d$ and the VP ratio of two incoherent Airy disks, as illustrated by the solid green line in Fig. 3(c). This criterion quantitatively delineates a limit that cannot be surpassed using incoherent light and is beneficial for discussing the separability of multispot patterns using coherent light.
@@
The area below the green line can be regarded as a super-resolution region, where the separation of adjacent spots is superior to the criterion for fixed spot spacing, or where the spot spacing is smaller than the criterion for a fixed VP ratio. According to the extended super-resolution criterion, the numerical VP values in Fig. 3(c) suggest the potential of FOI-designed holograms to generate super-resolution light spot arrays in the regime $d/r_{\text{A}}<1.8$. In particular, superior spot separation is anticipated for $1.71.3<d/r_{\text{A}}<$ 1.7. The RSS-designed holograms generate holograms surrounding the criterion but cannot overcome these limitations. The experimental VP ratios at $d/r_{\text{A}}=1.4$ in Fig. 3(c) exhibit moderately larger values than the numerical ones, suggesting the influence of extrinsic factors in the experiments, such as residual optical aberrations. Nevertheless, we emphasise that the results of the FOI-cost method still fall within the super-resolution regime. Refer to Supplementary Information for further details.
@@
As described in Methods, instabilities exist in the hologram design and optical reproduction, which can be distinguished by comparing the numerical and experimental results. A point of issue in this comparison is that errors in the numerical light utilisation efficiencies were similar for both the FOI and RSS methods, as listed in Table 1. Interestingly, the accuracy of the CGM procedure was the dominant contributor to these errors, which was the only common factor between the FOI and RSS methods. The CGM procedure contributes toward achieving higher numerical stability compared to self-consistent iterations like the IFTA method.
@@
In contrast, the numerical and experimental values of light utilisation efficiency are similar for the FOI method, implying that the FOI method is tolerant to experimental disturbances concerning light utilisation efficiency, which is a global property of light spot arrays. However, this argument does not apply to spot uniformity, which exemplifies the local properties of the spot arrays. The numerical results, including both the values and errors of spot uniformity, reveal distinct differences between the FOI and RSS methods, as outlined in Table 1. Thus, spot uniformity is affected by the selection of the cost function.
However, this difference reduced marginally in the experimental results, implying a nonnegligible contribution from experimental factors such as aberrations in optical trains. The FOI method generates super-resolution spot arrays of higher quality than the RSS method, and the structural stability of the FOI method is maintained to a certain extent even in the presence of experimental disturbances.
@@
The subsequent question is: how does the FOI method generate super-resolution spot arrays? A well-known example is the phase shift method [31] used in lithography. Although lithography differs from holography in terms of near and far-field light propagation phenomena, this technology has attempted to overcome the diffraction limits by controlling the phase distribution of coherent light. The RSS method can be classified under this type of approach owing to its high light utilisation efficiency and globally smooth CGHs, as illustrated in Fig. 1b.
@@
In contrast, the FOI-designed CGH shown in Fig. 1b presents a phase profile with prominent high-frequency components that cause diffusive light propagation. This leads to the characteristics of the FOI method, namely, low light utilisation efficiency. Therefore, FOI-designed CGHs may function as complex amplitude modulation phase holograms [32, 33, 34, 35] to remove the undesired light components that prevent super-resolution and cause interference, which induces instability in the holographically reproduced images. Conversely, the current results indicate a limitation of phase-only modulation approaches and the necessity of removing undesired light components to achieve super-resolution spot arrays.
@@
Before concluding the paper, we discuss the numerical properties of the FOI cost function. In CGH design problems, an “absolute” solution satisfying $I_{i}=T_{i}$ is not possible within a set of physical solutions at all instances due to the nature of light propagation. Consequently, CGHs designed using optimisation-based methods generally rely on cost functions. Here, it is essential to focus on the behaviour of the cost function rather than its superficial form. In fact, we can readily confirmed that $\sum(\tilde{I}_{i}-\tilde{T}_{i})^{2}=2-2\sum\tilde{I}_{i}\tilde{T}_{i}$ holds, indicating the equivalence of the FOI and RSS cost functions under Euclidean normalisation. Thus, one can infer that Euclidean normalisation itself plays a crucial role in generating CGHs for super-resolution spot array patterns. Although we were unable to establish a mathematically rigorous description of the significant changes in the reproduced images due to the cost functions, it is still worthwhile to discuss this briefly.
@@
Consider a test target pattern such that $T_{i}=c>0$ only at separated $M$ points, among $N$ lattice points in the image plane ($M\leq N$). The support for $T_{i}$ is denoted by $\alpha$ as follows: $\alpha=\{\forall i|T_{i}\neq 0\}$. The target pattern satisfies $T_{i}=1/M\ (i\in\alpha)$ for the total-sum normalisation, whereas $\tilde{T}_{i}=1/\sqrt{M}\ (i\in\alpha)$ holds for Euclidean normalisation. Therefore, the FOI cost function becomes $f_{\text{FOI}}=-(1/\sqrt{M})\sum_{i\in\alpha}\tilde{I}_{i}$. However, the RSS cost provides:
$\sum_{i}\left(I_{i}-T_{i}\right)^{2}$ $=2-{\sum_{i\neq k}}^{\prime}T_{i}T_{k}-2\sum_{i\in\alpha}I_{i}T_{i}-{\sum_{i\neq k}}^{\prime}I_{i}I_{k}$
$=\left(1+\frac{1}{M}\right)-\frac{2}{M}\sum_{i\in\alpha}I_{i}-{\sum_{i\neq k}}^{\prime}I_{i}I_{k},$ (5)
where $\sum I_{i}^{2}=1-\sum_{i\neq k}^{\prime}I_{i}I_{k}$ and a similar relation for $T_{i}$ was used for the aforementioned deformations. Note that the sum of the output values over $\alpha$ commonly works to reduce both FOI and RSS cost values, despite the difference in normalisation. However, an additional factor in $f_{\text{RSS}}$ reduces the cost value: the last term of Equation (5). This term accounts for the autocorrelation of the output pattern independent of the target, and it broadens the light spots, which is favourable for reducing $f_{\text{RSS}}$.
Conclusions
@@
This study discovered that the quality of narrow holographic optical lattice patterns could be significantly improved by utilising the FOI-cost method for hologram optimisation, as opposed to the traditional RSS-cost method. We have developed a super-resolution criterion specifically for arrays of light spots, which is based on the spacing of the spots and the VP ratio of the incoherently superimposed Airy disks. The multispot patterns generated in the present experiments satisfied this criterion, representing a key advancement for subsequent cold atom experiments employing holographic techniques [25]. The potential applications of this method are not limited to cold atoms; it also holds promise for laser processing [7, 8] and bioimaging [9]. In future, we intend to establish a comprehensive understanding of the relationship between other hologram design methods, such as IFTA and its derivatives, and optimisation-based methods. This exploration is expected to evolve, thereby expanding the potential application areas of the proposed method.

### v2

@@ IV.2 Spot uniformity and light utilization efficiency
Considering the scope of this study, the evaluation of the reproduced images should be based on the characteristic properties of light-spot arrays. One might assume that a value of FOI itself serves as a measure of the quality of the reproduced images, but the FOI reflects geometrical information of light-spot arrays only in an indirect manner. Therefore, we investigated the properties of the reproduced images in terms of the uniformity of the light-spot array and light utilization efficiency.
[…]
We also examined the effects of residual aberrations on the spot uniformity and light utilization efficiency. Recent literatures by one of the authors reported that even high-quality objective lenses cannot avoid residual aberrations of less than 0.02 $\lambda$ in root mean square [34, 35], which is far below detection limits of any existing methods. We calculated and analyzed the reproduced images by adding a typical aberration pattern observed for a plan-fluorite–type objective (Objective A in Ref. 35) to 20 FOI- and RSS-designed holograms. The results are shown in the middle row of Table 1, suggesting that aberrations can certainly be a factor degrading the spot uniformity. However, it is noted that the objective lens and optical train in this study are strictly different from those in Refs. 34 and 35, and that these results are for informational purposes.
@@
Figure 4: (a) Focal pattern (top) and profile (bottom) of Airy disk [NA = 0.75(air)]. Red line in the bottom plot shows cross-sectional profile corresponding to the focal pattern, where $r_{\text{A}}$ refers to the Airy-disk radius. (b) Focal patterns and profiles of incoherent sums of two Airy disks aligned at distances of $d=r_{\text{A}}$ (left) and $2r_{\text{A}}$ (right). In the bottom plots, thick green lines display the profiles of the incoherent sums, whereas thin red lines indicate those of individual Airy disks. Definitions of peak (P) and valley (V) are illustrated as well. (c) The relationship between the VP ratio and spot spacing for incoherent sums of adjacent Airy disks is depicted (solid green line), where the traditional Sparrow, Abbe, and Rayleigh criteria are plotted as purple crosses. Open symbols represent VP ratios for numerically reproduced images from FOI- (red open circles) and RSS-designed (blue open triangles) CGHs, while closed symbols represent those for the corresponding observed patterns, i.e., Figs. 3(c) (red closed circle) and (d) (blue closed triangle), respectively. Error bars (SEs) are determined similarly to those in Table 1 but are smaller than the size of the plot symbols, except for the experimental result of an RSS-designed CGH.
IV.3 Relation between spacing and separation of spot array
@@
FOI-designed holograms can generate high-quality spot arrays with extremely narrow spacing. To evaluate the behavior of spot arrays within the diffraction-limit regions, the focal patterns calculated under a high-NA [= 0.75 (air)] condition are displayed in Fig. 4. Note that all numerical results in Fig. 4 were calculated using vector Debye formulas [36, 37, 38, 39, 40], as the focal patterns of coherent light are affected by the vector nature of the light amplitude under NA conditions greater than 0.70. The Airy disk pattern calculated assuming circularly polarized light is illustrated in Fig. 4(a).
@@
The Rayleigh-limit conditions portrayed on the left in Fig. 4(b) present the adjacent light spots (Airy disks) as inadequately separated, while well-separated light spots, as displayed on the right of Fig. 4(b), are desirable for maintaining a high aspect ratio in applications such as laser processing and suppression of background noise in scanning laser microscopy. Therefore, a factor other than spot spacing is required for a proper evaluation of spot-array patterns while considering their applications.
@@
The valley-to-peak (VP) ratio of a spot array along with the inverse of the peak-to-valley ratio was introduced as a metric for spot separation. A smaller VP ratio indicates a clear separation of adjacent light spots. The VP ratio is advantageous because of its availability and robustness in derivation under unfavorable conditions compared to other metrics, such as the size of each light spot. To derive the VP ratio of the reproduced spot-array image, the peak positions were determined by fitting the reproduced images. The peak value refers to the mean of $I_{i}$ at the peak positions of all the bright spots, whereas the valley value refers to the mean of $I_{i}$ at all the midpoints between the adjacent peaks. Thus, the VP ratio was evaluated as the ratio of the valley value to the peak value.
@@
The blue (red) open symbols displayed in Fig. 4(c) represent the VP ratios of numerically reproduced images from hologram patterns designed by the FOI (RSS) method, whereas the blue (red) closed symbols indicate those derived from the experimental images. In terms of the VP ratio, the proposed FOI method outperforms the conventional RSS method in the range of $(1<d/r_{\text{A}}<2)$.
V Discussion
V.1 Extension of super-resolution criterion
@@
Now we attempt to extend the known super-resolution criteria to a definite, quantitative, and comprehensive one. In considering the relationship between physical quantities affecting spot separation, the traditional Rayleigh diffraction limit is established for an incoherent superposition of two Airy disks. Thus, this approach should be logically followed. We propose to determine whether a spot array exhibits “super-resolution” by examining the relationship between the spacing $d$ and the VP ratio of two incoherent Airy disks, as illustrated by the solid green line in Fig. 4(c). This criterion quantitatively delineates a limit that cannot be surpassed using incoherent light and is beneficial for discussing the separability of multispot patterns using coherent light.
@@
The area below the green line can be regarded as a super-resolution region, where the separation of adjacent spots is superior to the criterion for fixed spot spacing, or where the spot spacing is smaller than the criterion for a fixed VP ratio. According to the extended super-resolution criterion, the numerical VP values in Fig. 4(c) suggest the potential of FOI-designed CGHs to generate super-resolution light-spot arrays in the regime $d/r_{\text{A}}<1.8$. In particular, superior spot separation is anticipated for $1.3<d/r_{\text{A}}<1.7$. In this interval, the $d/r_{\text{A}}$-VP plot for the results of the RSS-designed CGHs lies around the extended criterion but not in the super-resolution region. On the other hand, in the region of $d/r_{\text{A}}>1.8$, the VP ratios of the FOI and RSS method do not differ significantly, and this trend continues for $d/r_{\text{A}}>2.0$, outside the range shown in Fig.  4(c). So, the RSS method will be preferable in terms of the light utilization efficiency under the condition of larger spot spacing.
The experimental VP ratios at $d/r_{\text{A}}=1.4$ in Fig. 4(c) exhibit moderately larger values than the numerical ones, suggesting the influence of extrinsic factors in the experiments, such as residual optical aberrations. Nevertheless, we emphasize that the results of the FOI method remain within the super-resolution regime. Refer to Supplemental Material [41] for further details.
So far, the discussion has been based on the Rayleigh criterion, but there are other super-resolution criteria, such as the Sparrow and Abbe criteria. Difference among those three criteria is reduced to the definition of separability bewteen two adjascent light spot; i.e., they differ only in the choice of the limit spacing at which the two Airy disks can be regarded as resolved. Our super-resolution criterion can be said to encompass all three of these, since all three criteria correspond to a single point (purple cross) on the solid green line in Fig. 4(c).
For example, the strictest Sparrow criterion, below which the valley between the two superimposed Airy disks disappears, can be expressed as a point corresponding to VP ratio of unity. At spot spacings closer than this, the valley value exceeds the peak, making the VP ratio as a physically meaningless measure. Thus, the Sparrow criterion also indicates the minimum spot spacing at which our super-resolution criterion is applicable.
Since this paper focuses on establishing and demonstrating the fundamentals of super-resolution criterion, the target pattern has been limited to square lattices. Therefore, quantitative discussion of non-square lattice patterns is not included here, but comments on applicability of our method to these patterns is provided in Appendix B.
V.2 Stability of CGHs against external disturbance
A sequence of pseudorandom numbers is completely determined by selecting an integer, known as the “random seed.” Different seeds can be selected to generate various hologram patterns for the same target patterns. The images reproduced from these different hologram patterns can be treated as independent statistical trials to evaluate the statistical errors of the physical quantities obtained through analyses.
In this study, the error of a physical quantity, defined as SE in a series of quantities, provides a benchmark for the stability of the reproduced images. It includes the effects of the non-uniqueness of the hologram design problem, the convergence criterion set for optimization in CGH design, the handling of phase values in CGHs (rounded to discrete values), and the numerical/experimental accuracy for hologram reproduction. The accuracy of the numerical holographic reproduction was equivalent to that of the FFT algorithm and is expected to be negligibly small (less than $5\times 10^{-5}$ relative to the root-mean-square error, even for single-precision FFT, as noted in https://pypi.org/project/pyvkfft/). This fact justifies neglecting the errors of numerically derived quantities. Therefore, the instabilities in the hologram design and optical reproduction stages, namely, the intrinsic and extrinsic instabilities, can be distinguished by comparing the numerical and experimental results presented in Table 1.
@@ V.2 Stability of CGHs against external disturbance
As the first point of issue in Table 1, errors in the numerical light utilization efficiencies were similar for both the FOI and RSS methods. This observation suggests that the accuracy of the CGM procedure, the only common factor between the FOI and RSS methods, can be assigned as the dominant contributor to these errors. From a different point of view, the CGM procedure contributes toward achieving higher numerical stability compared to self-consistent iterations like the IFTA method.
@@
In contrast, the numerical and experimental values of light utilization efficiency are similar for the FOI method. This result implies that the FOI method is tolerant to experimental disturbances concerning light utilization efficiency, which is a global property of light-spot arrays. However, the argument does not apply to spot uniformity, which exemplifies the local properties of the spot arrays. The numerical results, including both the values and errors of spot uniformity, reveal distinct differences between the FOI and RSS methods, as outlined in Table 1. Thus, spot uniformity is affected by the selection of the cost function.
In practical optical systems, multiple factors such as lens aberrations, misalignment, vignetting, the uniformity and phase flatness of incident light, and phase-modulation characteristics of a SLM device interact in complicated ways to degrade the quality of the reproduced images. Since it is almost impossible to evaluate all these factors via numerical simulation, we restricted our attention to effects of aberrations on the reproduced images. Numerical results in Table 1 show that the aberrations certainly affect significanlty the spot uniformity but do slightly the light utilization efficiency. It is noted again that the numerical evaluation of the aberrations is performed for informational purposes and not the subject of quantitative comparison. Nevertheless, it is meaningful to highlight the effects of a tiny phase distortion expected in practical optical systems.
The remaining question on the results of in Table 1 is why there is a significant discrepancy between numerical and experimental values of the light utilization efficiency for the RSS method. One possible qualitative explanation is as follows: In FOI-designed holograms, where diffusive components dominate, effects of extrinsic disturbance in experimental systems can be buried in the scattered elements that contribute little to holographic output images. As a result, the FOI method presents superficial tolelance to extrinsic disturbance in practical optical systems. In contrast, in RSS-designed holograms, where majority parts of light components contribute to the output images reproduced as results of strong interference with each other, even the minor disturbance can cause significant impact on the output images. In relation with the above explanation, function of FOI-designed holograms is analyzed in the next subsection.
V.3 Physical interpretation of FOI-designed CGHs
@@
The subsequent question is: how does the FOI method generate super-resolution spot arrays? A well-known example is the phase shift method [42] used in lithography. Although lithography differs from holography in terms of near and far-field light propagation phenomena, this technology has attempted to overcome the diffraction limits by controlling the phase distribution of coherent light. The RSS method can be classified under this type of approach owing to its high light utilization efficiency and globally smooth CGHs, as illustrated in Fig. 1(b).
@@
In contrast, the FOI-designed CGH shown in Fig. 1(b) presents a phase profile with prominent high-frequency components that cause diffusive light propagation. This leads to the characteristics of the FOI method, namely, low light utilization efficiency. Therefore, FOI-designed CGHs may function as complex amplitude modulation phase holograms [43, 44, 45, 46] to remove the undesired light components that prevent super-resolution and cause interference, which induces instability in the holographically reproduced images. Conversely, the current results indicate a limitation of phase-only modulation approaches and the necessity of removing undesired light components to achieve super-resolution spot arrays.
V.4 Properties of FOI cost function
@@
Before concluding the paper, we discuss the numerical properties of the FOI cost function. In CGH design problems, an “absolute” solution satisfying $I_{i}=T_{i}$ is not possible within a set of physical solutions at all instances due to the nature of light propagation. Consequently, CGHs designed using optimization-based methods generally rely on cost functions. Here, it is essential to focus on the behavior of the cost function rather than its superficial form. In fact, we can readily confirmed that $\sum(\tilde{I}_{i}-\tilde{T}_{i})^{2}=2-2\sum\tilde{I}_{i}\tilde{T}_{i}$ holds, indicating the equivalence of the FOI and RSS cost functions under Euclidean normalization. Thus, one can infer that Euclidean normalization itself plays a crucial role in generating CGHs for super-resolution spot array patterns. Although we were unable to establish a mathematically rigorous description of the significant changes in the reproduced images due to the cost functions, it is still worthwhile to discuss this briefly.
@@
Consider a test target pattern such that $T_{i}=c>0$ only at separated $M$ points, among $N$ lattice points in the image plane ($M\leq N$). The support for $T_{i}$ is denoted by $\alpha$ as follows: $\alpha=\{\forall i|T_{i}\neq 0\}$. The target pattern satisfies $T_{i}=1/M\ (i\in\alpha)$ for the total-sum normalization, whereas $\tilde{T}_{i}=1/\sqrt{M}\ (i\in\alpha)$ holds for Euclidean normalization. Therefore, the FOI cost function becomes $f_{\text{FOI}}=-(1/\sqrt{M})\sum_{i\in\alpha}\tilde{I}_{i}$. However, the RSS cost provides: {align} ∑_i ( I_i - T_i )^2 = 2 - ∑_i ≠k^′T_i T_k - 2 ∑_i ∈α I_i T_i - $i I_k \notag$
[…]
VI Conclusions
@@
This study discovered that the quality of narrow holographic optical lattice patterns could be significantly improved by utilizing the FOI method for optimization-based hologram design, as opposed to the traditional RSS method. We have developed a super-resolution criterion specifically for arrays of light spots, which is based on the spacing of the spots and the VP ratio of the incoherently superimposed Airy disks. The multispot patterns generated in the present experiments satisfied this criterion, representing a key advancement for subsequent cold atom experiments employing holographic techniques [31, 14]. The potential applications of this method are not limited to cold atoms; it also holds promise for laser processing [7, 8] and bioimaging [9]. Our method enables flexible and high-precision spot alignment without modifying the optical components used in the system, making it a versatile technique for a wide range of applications. When combined with a high-resolution optical system, our approach has the potential to generate extremely fine patterns. In future, we intend to establish a comprehensive understanding of the relationship between other hologram design methods, such as IFTA and its derivatives, and optimization-based methods. This exploration is expected to evolve, thereby expanding the potential application areas of the proposed method.
Acknowledgements.
The authors are grateful to H. Toyoda, T. Inoue, and Y. Ohtake of HPK for their encouragement throughout this work, and to T. Watanabe for kind supports on fabrication of optics. We also appreciate discussion and kind support provided by Prof. K. Ohmori of IMS as the leader of ultrafast cold-atom quantum simulator projects. Part of this study was performed under support of JSPS Grant-in-Aid for Specially Promoted Research (grant no. 16H06289). T.T. and S.d.L. also acknowledge partial support by JSPS Grant-in-Aid for Research Activity Start-up (grant nos. 19K23431 and 19K23429, respectively) as well as by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) JPMXS0118069021 and JST Moonshot R&D Program Grant Number JPMJMS2269.
Appendix A Incident light conditions and reproduced image quality
In this section, we present the results of numerical simulations on how amplitude distribution of incident light affects the quality of holographically reproduced images. Although holograms were designed by assuming a top-hat beam as incident light, i.e., uniform illumination over a circular region around the center of hologram plane, it is difficult in practice to generate precise top-hat beams. For this reason, a truncated Gaussian beam is often adopted in practical experiments as a pseudo top-hat beams. Therefore, we numerically simulated the spot uniformity, light utilization efficiencies and VP ratio when the holograms are designed with a top-hat beam as incident light but reproduced by using a truncated Gaussian beam. The simulation conditions not mentioned below are set to be the same as those in Sec. III.
We evaluated these properties from hologram images, which are reproduced numerically under illumination of truncated Gaussian beams, for 20 FOI-holograms designed in Sec. III. Illumination profiles of the truncated Gaussian beams were determied by extracting the central area of 4.5 mm radius (corresponding to 360 px on the active area of the SLM) of Gaussian beams with $1/e^{2}$ radius of 5, 10, and 20 mm. The results are summarized in Table 2 together with the values obtained from numerical results reproduced under illumiation of an ideal top-hat beam for reference. In the case of 20 mm radius, which is close to the experimental conditions, there is no significant change in any of the properties, which suggests that it is reasonable to approximate the incident beam as a top-hat beam. On the other hand, if incident light is narrowed down to a radius of 5 mm, the quality degradation of the reproduced image becomes notable.
Table 2: Spot uniformity, light utilization efficiency, and VP ratio of light-spot arrays that are numerically reproduced from FOI-designed CGHs illuminated by top-hat beam and truncated Gaussian beams of various radii.
$1/e^{2}$ radius Spot uniformity Util. eff. VP ratio
(Top-hat) $1.21(4)\times 10^{-2}$ 0.120(3) $1.00(5)\times 10^{-2}$
20 mm $1.41(4)\times 10^{-2}$ 0.116(3) $1.00(4)\times 10^{-2}$
10 mm $2.6(1)\times 10^{-2}$ 0.108(3) $1.10(4)\times 10^{-2}$
5 mm $9.9(3)\times 10^{-2}$ 0.076(2) $1.67(5)\times 10^{-2}$
Appendix B Applicability of our method to non-square lattice patterns
For the reasons explained in Sec. V.1, the experiments in this study were limited to square lattice patterns as the target pattern. However, to fully leverage the advantages of holographic techniques, it is worth discussing the applicability to other lattice patterns. Our super-resolution criterion for multispot patterns is defined based on the average VP ratio between adjacent spots and is therefore applicable to non-square lattice patterns. Figure 5 shows numerically reproduced images of hexagonal, kagome, and triangular lattices generated by the FOI method, with spot spacings of 1.9, 1.9, and 2.3 pixels on the image plane, respectively. Compared to square lattices, these alternative lattice patterns tend to allow for greater spacing between spots while still maintaining clear separation. In the diffraction-limited region, multispot patterns behave not as collections of independent spots but rather as spatially distributed patterns like higher-order transversal light modes sometimes exhibiting non-circular spot shapes depending on the geometry of each lattice. Although investigating the reasons behind these observations is beyond the scope of this study, one plausible hypothesis is that the square lattice pattern is relatively close to physically realizable Hermite-Gaussian modes, which contributes to its stability.
Figure 5: Reproduced images of FOI-holograms with hexagonal, kagome, and triangular lattices as target patterns. The spot intervals are 1.9, 1.9, and 2.3 pixels, respectively.